OpenStudy (itsmeyourfren):

PLEASE HELPP :) Multiply (2.4 ⋅ 1014) ⋅ (4 ⋅ 107). Express the answer in scientific notation. 9.6 ⋅ 1021 9.6 ⋅ 1022 96 ⋅ 1021 96 ⋅ 1022

jhonyy9 (jhonyy9):

ok but i think there are 10^14 and 10^7 yes ?

jhonyy9 (jhonyy9):

for scientific notation just so in this way make sens

jhonyy9 (jhonyy9):

so in this case this is easy 10^14 *10^7 = 10^21and 2,4*4 = 9,6 hope helped you

jhonyy9 (jhonyy9):

the first choice is correct,right sure